1) Hydrogen peroxide, H2O2(aq), decomposes to H2O(l) and O2(g) in a reaction that is first order in H2O2 and has a rate constant k = 1.06×10−3 min−1 at a given temperature. How long will it take for 15% of a sample of H2O2 to decompose?

2)The decomposition of nitrogen dioxide, NO2, into nitrogen monoxide and oxygen at a high temperature is second-order in NO2. The rate constant for this reaction is 3.40 L/mol×min. Determine the time needed for the concentration of NO2 to decrease from 2.00 mol/L to 1.50 mol/L.   
Answer #1

1)
we have:
[H2O2]o = 100 M (Let initial concentration be 100 M)
15 % has decomposed
SO, remaining is 85 %
[H2O2] = 85 M
k = 1.06*10^-3 min-1

use integrated rate law for 1st order reaction
ln[H2O2] = ln[H2O2]o - k*t
ln(85) = ln(100) - 1.06*10^-3*t
4.4427 = 4.6052 - 1.06*10^-3*t
1.06*10^-3*t = 0.1625
t = 153 min
Answer: 153 min

Only 1 question at a time please